# Multigranular Evaluation for Brain Visual Decoding

### v1

@@ Appendix D Qualitative Comparative Analysis
This section presents additional qualitative results and analyses, including brain visual decoding reconstructions (Sec. D.1), detailed captions for both reconstructions and visual stimuli (Sec. D.2), parsed caption annotations (Sec. D.3), and ablations on using different MLLMs and prompts (Sec. D.4 and Sec. D.5). The results are primarily based on methods applied to NSD allen2022massive for a comprehensive and in-depth examination.
D.1 Brain Decoding Reconstruction
Fig. S2 presents qualitative comparisons of brain visual decoding methods on the NSD dataset, including the original visual stimuli and reconstructions from SDRecon takagi2023improving , BrainDiffuser ozcelik2023brain , MindEye scotti2023reconstructing , DREAM xia2024dream , MindEye2 scotti2024mindeye2 , MindBridge wang2024mindbridge , UMBRAE xia2024umbrae , NeuroPictor huo2024neuropictor , NeuroVLA shen2024neuro , UniBrain wang2024unibrain , STTM liu2025see , MindTuner gong2025mindtuner , and BrainGuard tian2025brainguard . These results encompass the majority of current brain visual decoding methods on NSD, providing a representative and comprehensive overview that reflects the latest advancements and trends in the field.
@@ D.1 Brain Decoding Reconstruction
D.2 Detailed Captioning Comparison
@@
We highlight the objects, attributes, and relationships extracted from the captions using the pipeline introduced in the main paper. These provide a structured and interpretable representation of the decoded contents. It should be noted that the annotations in the original captions shown below are intended for intuitive understanding and differ slightly from the parsed captions used for final evaluation, which are stored as structured data. The actual parsed results are in Sec. D.3.
@@
    The image captures a bustling scene at an terminal. The terminal is characterized by a curved ceiling, adorned with a pattern of diagonal lines on the ceiling that add a sense of depth and dimension to the space. A large sign, emblazoned with the words on the sign—"FLIGHTS" and "GATE"—hang from the ceiling, guiding travelers to their respective destinations. The terminal is abuzz with activity. Numerous people walking can be seen moving around, each engrossed in their own journey. Their movements create a dynamic tableau, add life and add energy to the otherwise static environment. The terminal is also home to a variety of scattered vehicles, including cars and buses, which are in the terminal alongside the people. These vehicles, along with the people, contribute to the overall sense of movement and activity in the terminal. Despite the monochrome palette of the image, the scene is rich in detail and texture. The interplay of light and shadow, the intricate patterns on the ceiling, and the varied elements within the terminal all come together to create a vibrant, engaging image. The image is a testament to the constant motion and energy that characterizes life at an airport terminal.
@@ D.3 Parsed Descriptions
D.4 Captioning with Different MLLMs
The following two subsections present ablation examples evaluating the impact of different MLLMs and prompting strategies on detailed captioning. For clarity and brevity, we report only the parsed descriptions of objectsattributes, and relationships in the structured format. Despite variations in the generated captions across different MLLMs and prompts, the evaluation scores and method ranking remain consistent, as shown in Fig. 3, with more detailed scores in the quantitative evaluation (see  Tab. S9 and Tab. S10 in Sec. E.2).
@@ D.4 Captioning with Different MLLMs
D.5 Captioning with Different Prompts
@@
D.6 Comparison of Multigranular Segmentation Results
Fig. S3 shows the comparison of multigranular segmentation results between reference images and reconstructions. The segmentations are visualized at multiple levels, including semantic, instance, and part granularity. Unlike previous metrics that evaluate only low-level consistency with the reference using a single black-box score, our segmentation-based multigranular metric provides more interpretable insights. IoU and AP scores are higher when the reconstructions are structurally aligned with the reference. Taking examples in Fig. S3, reconstructions where an airplane or zebra appears in the correct position and orientation receive higher scores—a property that holds across different granularities. For instance, BrainDiffuser ozcelik2023brain yields lower scores due to incorrect airplane orientation, while MindEye scotti2023reconstructing experiences score drops resulting from perspective shifts caused by camera zoom. Overall, most methods produce reconstructions that preserve the high-level semantic structure of the original images, as reflected in the alignment of major object categories. However, finer-grained details, such as part boundaries or object instances in crowded scenes, are often less accurately recovered, suggesting limitations in capturing precise spatial layouts and textures.
@@ Appendix E Quantitative Comparative Analysis
E.1 Evaluating NSD with the BASIC
This section presents a performance comparison of visual decoding methods on NSD allen2022massive using the proposed BASIC metrics, which capture both structural (BASIC-L) and semantic (BASIC-H) alignment. The final BASIC score is computed as the average of BASIC-L and BASIC-H, reflecting the equal importance of low-level structural and high-level semantic alignment in brain visual decoding evaluation. Fig. S4 presents the results, with the values of the top-5 best performing methods annotated for each dimension. For structural alignment NeuroPictor huo2024neuropictor , STTM liu2025see , MindEye2 scotti2024mindeye2 , BrainGuard tian2025brainguard , and DREAM xia2024dream achieve the highest BASIC-L scores, reflecting better spatial reconstruction. In contrast, semantic alignment (BASIC-H) is best captured by NeuroVLA shen2024neuro , DREAM xia2024dream , STTM liu2025see , BrainGuard tian2025brainguard , and MindTuner gong2025mindtuner , indicating superior semantic preservation. NeuroPictor huo2024neuropictor ranks highest in the composite BASIC score, followed by STTM liu2025see , BrainGuard tian2025brainguard , MindEye2 scotti2024mindeye2 , and DREAM xia2024dream , reflecting more balanced performance across both perceptual dimensions. While NeuroVLA shen2024neuro excels in semantic similarity, it underperforms in structural precision, suggesting a trade-off between high-level conceptual encoding and spatial accuracy. These results demonstrate the complementary nature of BASIC-L and BASIC-H and underscore the importance of multigranular evaluation for benchmarking brain visual decoding.
@@ E.1 Evaluating NSD with the BASIC
E.2 Stability across Evaluation Configurations
@@
Tab. S9 presents the quantitative results obtained using different captioning prompts. Three prompt variants, same as in Sec. D.5, are considered: (1) “Describe the image in detail.”, (2) “Describe the image in detail, focusing on visible objects.”, and (3) “Describe the image in detail, focusing on visible objects and their relationships.” Despite slight differences in phrasing, the evaluation scores remain consistent across prompts, indicating that our metric is robust to moderate prompt variations and does not overly rely on specific wording. The MLLM used here is LLaVA-1.6-7B liu2023visual . The metric remains consistent and discriminative even with very close scores (refer to the demonstrated scores of UMBRAE xia2024umbrae , NeuroPictor huo2024neuropictor , MindEye scotti2023reconstructing , and MindEye2 scotti2024mindeye2 ).
Tab. S10 presents the quantitative evaluation results obtained using different MLLMs, including LLaVA-1.5-7B, LLaVA-1.5-13B, LLaVA-1.6-7B, LLaVA-1.6-13B, LLaVA-1.6-34B, and GPT-4o. Despite variations in model version and scale, the overall metric scores remain stable, and the relative method rankings are mostly preserved. This consistency across MLLMs demonstrates the robustness of our evaluation framework with respect to the choice of captioning model.
Tab. S11 presents a cross-evaluation, where reconstruction captions from each method are generated using one MLLM and assessed against captions produced by different MLLMs as references. Since no human-ranked scores are available for relevance analysis (i.e., consistency with human expert judgments using metrics such as Pearson correlation, coefficient of determination, or Kendall rank correlation), and obtaining reliable human-annotated relative rankings for each method is challenging, we instead adopt a cross-model evaluation strategy. Specifically, we assess performance by varying the MLLM models used as the producer (i.e., generating captions for the reconstructed images) and the evaluator (generating reference captions). This cross-model setup reveals consistency across MLLMs, suggesting that reliance on human-written “ground truth” captions is not strictly necessary for evaluation. Tab. S12 presents the detailed subscores for BrainDiffuser in the cross-model evaluation.
Tab. S13 summarizes the effect of varying box and text thresholds on the multigranular segmentation evaluation. While threshold changes influence the individual sub-scores, the relative rankings of methods remain consistent within each configuration. The final BASIC-L score demonstrates notable stability across all threshold settings, despite fluctuations in component metrics.
@@
Table S11: Cross-model evaluation. Each method’s reconstruction captions (recon) are generated using one MLLM and evaluated against reference captions (stimuli) produced by different MLLMs. This setup assesses the impact of using various MLLMs as the caption producer and the evaluator.
@@ Appendix F Limitation and Broader Impact
Despite our efforts in developing a versatile and robust evaluation pipeline, there are still challenges in capturing nuances in brain visual decoding, such as exploring additional cognitive dimensions like emotional or attentional states, to enable a more holistic evaluation. This limitation stems from current datasets lacking sufficient diversity in cognitive states or subject demographics. Furthermore, our method targets recent decoding techniques that can generate clear and semantically relevant images. The aspect of image quality is therefore not considered a key criterion. This makes the framework less suitable for assessing early-stage visual coding work that often contains artifacts, blurriness, or difficult-to-recognize semantics. The reliance on large-scale multimodal language models also presents challenges in terms of computational cost and susceptibility to hallucination.
@@
This presented evaluation framework provides a more detailed, interpretable, and standardized method for assessing visual decoding models. These enhanced evaluation capabilities facilitates a deeper understanding of how models interpret and reconstruct visual information, which is crucial for advancing brain-inspired models, brain-computer interfaces, and assistive technologies. To address privacy and security concerns, particularly those surrounding neurodata and potential technological misuse, it is essential to rigorously assess the ethical implications of model deployment and implement robust safeguards for participant data before applying our evaluation framework.

### v2

@@ Appendix D Qualitative Comparative Analysis
This section presents additional qualitative results and analyses, including brain visual decoding reconstructions, detailed captions for both reconstructions and stimuli, parsed caption annotations, and ablations on using different MLLMs and prompts. The results are primarily based on methods applied to NSD (Allen et al. 2022) for a comprehensive and in-depth examination.
Brain Decoding Reconstruction
Fig. S2 presents qualitative comparisons of brain visual decoding methods on the NSD dataset, including the original visual stimuli and reconstructions from SDRecon (Takagi and Nishimoto 2023b), BrainDiffuser (Ozcelik and VanRullen 2023), MindEye (Scotti et al. 2023), DREAM (Xia et al. 2024b), MindEye2 (Scotti et al. 2024), MindBridge (Wang et al. 2024a), UMBRAE (Xia et al. 2024a), NeuroPictor (Huo et al. 2024), NeuroVLA (Shen et al. 2024), UniBrain (Wang et al. 2024b), STTM (Liu et al. 2025), MindTuner (Gong et al. 2025), and BrainGuard (Tian et al. 2025). These results encompass the majority of current brain visual decoding methods on NSD, providing a representative and comprehensive overview that reflects the latest advancements and trends in the field.
@@
Detailed Captioning Comparison
@@
We highlight the objects, attributes, and relationships extracted from the captions using the pipeline introduced in the main paper. These provide a structured and interpretable representation of the decoded contents. It should be noted that the annotations in the original captions shown below are intended for intuitive understanding and differ slightly from the parsed captions used for final evaluation, which are stored as structured data. The actual parsed results are below.
@@
    The image captures a bustling scene at an terminal. The terminal is characterized by a curved ceiling, adorned with a pattern of diagonal lines on the ceiling that add a sense of depth and dimension to the space. A large sign, emblazoned with the words on the sign—”FLIGHTS” and ”GATE”—hang from the ceiling, guiding travelers to their respective destinations. The terminal is abuzz with activity. Numerous people walking can be seen moving around, each engrossed in their own journey. Their movements create a dynamic tableau, add life and add energy to the otherwise static environment. The terminal is also home to a variety of scattered vehicles, including cars and buses, which are in the terminal alongside the people. These vehicles, along with the people, contribute to the overall sense of movement and activity in the terminal. Despite the monochrome palette of the image, the scene is rich in detail and texture. The interplay of light and shadow, the intricate patterns on the ceiling, and the varied elements within the terminal all come together to create a vibrant, engaging image. The image is a testament to the constant motion and energy that characterizes life at an airport terminal.
@@
Captioning with Different MLLMs
The following two subsections present ablation examples evaluating the impact of different MLLMs and prompting strategies on detailed captioning. For clarity and brevity, we report only the parsed descriptions of objectsattributes, and relationships in the structured format. Despite variations in the generated captions across different MLLMs and prompts, the evaluation scores and method ranking remain consistent, as shown in Fig. 3, with more detailed scores in the quantitative evaluation (see  Table S9 and Table S10 ).
@@
Captioning with Different Prompts
@@
Comparison of Multigranular Segmentation Results
Fig. S3 shows the comparison of multigranular segmentation results between reference images and reconstructions. The segmentations are visualized at multiple levels, including semantic, instance, and part granularity. Unlike previous metrics that evaluate only low-level consistency with the reference using a single black-box score, our segmentation-based multigranular metric provides more interpretable insights. IoU and AP scores are higher when the reconstructions are structurally aligned with the reference. Taking examples in Fig. S3, reconstructions where an airplane or zebra appears in the correct position and orientation receive higher scores—a property that holds across different granularities. For instance, BrainDiffuser (Ozcelik and VanRullen 2023) yields lower scores due to incorrect airplane orientation, while MindEye (Scotti et al. 2023) experiences score drops resulting from perspective shifts caused by camera zoom. Overall, most methods produce reconstructions that preserve the high-level semantic structure of the original images, as reflected in the alignment of major object categories. However, finer-grained details, such as part boundaries or object instances in crowded scenes, are often less accurately recovered, suggesting limitations in capturing precise spatial layouts and textures.
@@ Appendix E Quantitative Comparative Analysis
Evaluating NSD with the BASIC
This section presents a performance comparison of visual decoding methods on NSD (Allen et al. 2022) using the proposed BASIC metrics, which capture both structural (BASIC-L) and semantic (BASIC-H) alignment. The final BASIC score is computed as the average of BASIC-L and BASIC-H, reflecting the equal importance of low-level structural and high-level semantic alignment in brain visual decoding evaluation. Fig. S4 presents the results, with the values of the top-5 best performing methods annotated for each dimension. For structural alignment NeuroPictor (Huo et al. 2024), STTM (Liu et al. 2025), MindEye2 (Scotti et al. 2024), BrainGuard (Tian et al. 2025), and DREAM (Xia et al. 2024b) achieve the highest BASIC-L scores, reflecting better spatial reconstruction. In contrast, semantic alignment (BASIC-H) is best captured by NeuroVLA (Shen et al. 2024), DREAM (Xia et al. 2024b), STTM (Liu et al. 2025), BrainGuard (Tian et al. 2025), and MindTuner (Gong et al. 2025), indicating superior semantic preservation. NeuroPictor (Huo et al. 2024) ranks highest in the composite BASIC score, followed by STTM (Liu et al. 2025), BrainGuard (Tian et al. 2025), MindEye2 (Scotti et al. 2024), and DREAM (Xia et al. 2024b), reflecting more balanced performance across both perceptual dimensions. While NeuroVLA (Shen et al. 2024) excels in semantic similarity, it underperforms in structural precision, suggesting a trade-off between high-level conceptual encoding and spatial accuracy. These results demonstrate the complementary nature of BASIC-L and BASIC-H and underscore the importance of multigranular evaluation for benchmarking brain visual decoding.
@@
Stability across Evaluation Configurations
@@
Table S9 presents the quantitative results obtained using different captioning prompts. Three prompt variants, same as above are considered: (1) “Describe the image in detail.”, (2) “Describe the image in detail, focusing on visible objects.”, and (3) “Describe the image in detail, focusing on visible objects and their relationships.” Despite slight differences in phrasing, the evaluation scores remain consistent across prompts, indicating that our metric is robust to moderate prompt variations and does not overly rely on specific wording. The MLLM used here is LLaVA-1.6-7B (Liu et al. 2023). The metric remains consistent and discriminative even with very close scores (refer to the demonstrated scores of UMBRAE (Xia et al. 2024a), NeuroPictor (Huo et al. 2024), MindEye (Scotti et al. 2023), and MindEye2 (Scotti et al. 2024)).
Table S10 presents the quantitative evaluation results obtained using different MLLMs, including LLaVA-1.5-7B, LLaVA-1.5-13B, LLaVA-1.6-7B, LLaVA-1.6-13B, LLaVA-1.6-34B, and GPT-4o. Despite variations in model version and scale, the overall metric scores remain stable, and the relative method rankings are mostly preserved. This consistency across MLLMs demonstrates the robustness of our evaluation framework with respect to the choice of captioning model.
Table S11 presents a cross-evaluation, where reconstruction captions from each method are generated using one MLLM and assessed against captions produced by different MLLMs as references. Since no human-ranked scores are available for relevance analysis (i.e., consistency with human expert judgments using metrics such as Pearson correlation, coefficient of determination, or Kendall rank correlation), and obtaining reliable human-annotated relative rankings for each method is challenging, we instead adopt a cross-model evaluation strategy. Specifically, we assess performance by varying the MLLM models used as the producer (i.e., generating captions for the reconstructed images) and the evaluator (generating reference captions). This cross-model setup reveals consistency across MLLMs, suggesting that reliance on human-written “ground truth” captions is not strictly necessary for evaluation. Table S12 presents the detailed subscores for BrainDiffuser in the cross-model evaluation.
Table S13 summarizes the effect of varying box and text thresholds on the multigranular segmentation evaluation. While threshold changes influence individual sub-scores, the relative rankings of methods remain consistent within each configuration. The final BASIC-L score demonstrates notable stability across all threshold settings, despite fluctuations in component metrics.
@@
Table S11: Cross-model evaluation. Each method’s reconstruction captions (recon) are generated using one MLLM and evaluated against reference captions (stimuli) produced by another different MLLMs. This setup assesses the impact of using various MLLMs as the caption producer and the evaluator.
@@ Appendix F Limitation and Broader Impact
Despite our efforts in developing a versatile and robust evaluation pipeline, there are still challenges in capturing nuances in brain visual decoding, such as exploring additional cognitive dimensions like emotional or attentional states, to enable a more holistic evaluation. This limitation stems from current datasets lacking sufficient diversity in cognitive states or subject demographics. Furthermore, our method targets recent decoding techniques that can generate clear and semantically relevant images. The aspect of image quality is therefore not considered a key criterion. This makes the framework less suitable for assessing early-stage visual coding work that often contains artifacts, blurriness, or difficult-to-recognize semantics. The reliance on large-scale MLLMs also presents challenges in terms of computational cost and susceptibility to hallucination.
@@
This presented framework provides a more detailed, interpretable, and standardized method for assessing visual decoding models. These enhanced evaluation capabilities facilitates a deeper understanding of how models interpret and reconstruct visual information, which is crucial for advancing brain-inspired models, brain-computer interfaces, and assistive technologies. To address privacy and security concerns, particularly those surrounding neurodata and potential misuse, it is essential to rigorously assess ethical implications and implement safeguards for participant data before applying the evaluation.